Write the following system of equations in the form AX = B, and calculate the solution using the equation X=A^{-1}B

x + y = -6

3x - y = -2
Answer #1

x+y = -6 32 -8 -2 writter as can be The systein of equations A X = B where A= 3 6. B V AX=B - A B A = det A -3 so X A B 4 - -